A sump pump (used to drain water from the basement of houses built below the water table) is draining a flooded basement at the rate of 0.850 L/s, with an output pressure of 4.00 ✕ 105 N/m2. You may neglect frictional losses in both parts of this problem. (a) The water enters a hose with a 3.00 cm inside diameter and rises 2.90 m above the pump. What is its pressure (in N/m2) at this point? (b) The hose goes over the foundation wall, losing 1.20 m in height, and widens to 4.00 cm diameter. What is the pressure now (in N/m2)?
Answer #1

_____________________________________________________

a)

P = P1 - P2

= (4.00 * 105) - (1000 * 9.8 * 2.9)

= 3.72 * 105 N/m2

_____________________________________________________

b)

V = v1 A1

v1 = 0.850 * 10-3 / pi * 0.0152 = 1.20 m/s

v2 = 0.850 * 10-3 / pi * 0.022 = 0.676 m/s

P2 = P1 + 1/2 p (v12 - v22) + p g (h1 - h2)

= (4.00 * 105) + 1/2 * 1000 (1.202 - 0.6762) + 1000 * 9.8 (0 - (2.9 - 1.2))

= 3.84 * 105 N/m2

_____________________________________________________

Please rate
if any mistake in this answer please comment i will clarify your doubt . thank you